# On the Effectiveness of Methods and Metrics for Explainable AI in Remote Sensing Image Scene Classification

### v1

@@ III C 1 Background
This measures how many of the $K$ most important features, according to $\bm{e}$, match the pixel‐level reference set $S_{gt}$. If all top-$K$ pixels are within the reference mask, then $\Psi_{\mathrm{TKI}}=1$, therefore a higher score is better.
@@
The RRA [31] measures the concentration of high-intensity relevance values within the pixel-level reference map. Let $|S_{\text{tot}}|$ be the size of the pixel-level reference map mask. Select the $K$ highest relevance values, according to $\Phi(\bm{x})$, count how many of these values lie within the actual pixel locations, and divide by $|S_{\text{tot}}|$. This can be expressed as:
$P_{\text{top }K}=\left\{p_{1},p_{2},\ldots,p_{K}\mid R_{p_{1}}>R_{p_{2}}>\ldots>R_{p_{K}}\right\},$ (33)
where $P_{\text{top }K}$ is the set of pixels with relevance values $R_{p_{1}},R_{p_{2}},\ldots,R_{p_{K}}$ sorted in decreasing order until the $K$-th pixel. The RRA metric $\Psi_{\mathrm{RRA}}$ is then computed as:
$\Psi_{\mathrm{RRA}}=\frac{\left|P_{\text{top }K}\cap S_{\text{tot}}\right|}{|S_{\text{tot}}|}.$ (34)
III C 2 Limitations
@@
In RS scene classification, label dependencies arise when the presence or absence of one label influences the likelihood of another, which is a common scenario in RS where certain objects (e.g., ships) typically appear in specific environments (e.g., sea). These dependencies are not considered by localization metrics. Furthermore, the reliability of localization metrics heavily depends on accurate and up-to-date pixel-level reference maps. Obtaining such annotations is costly and prone to errors. Inaccuracies or obsolescence in the reference maps can compromise the localization reliability, potentially leading to incorrect assessments. Additionally, most localization metrics become unreliable, when the target class spans a large area of the input image. Let us consider the TKI metric in the scenario where we have a single label for the input image. Here, the binary reference mask $\mathbf{s}_{gt}$ for the target label becomes a matrix of ones, encompassing all pixels. Thus, $\Psi_{\mathrm{TK}}$ inherently achieves its maximum value of 1, regardless of the explanation $\bm{e}$. Since $\mathbf{s}_{gt}$ includes every pixel, the intersection $\mathbf{s}_{1-K}=\mathbf{r}_{1-K}\cap\mathbf{s}_{gt}$ reduces to $\mathbf{r}_{1-K}$ itself, as all top-ranked indices in $\mathbf{r}_{1-K}$ necessarily lie within $\mathbf{s}_{gt}$. Consequently, the numerator and denominator of $\Psi_{\mathrm{TK}}$ both equal $K$, yielding $\Psi_{\mathrm{TK}}=1$. In this case, the metric cannot differentiate meaningful explanations from arbitrary ones. This phenomenon is well-documented for simpler localization metrics [55], such as the Pointing Game [56]. However, in RS scene classification this also creates challenges for more sophisticated metrics, as in most RS scene classification tasks, as classes tend to span larger areas.
III D Complexity
III D 1 Background
@@
The fourth category we consider is complexity, which quantifies the simplicity and interpretability of explanations, ensuring that they are useful for human understanding. These metrics evaluate properties such as sparsity, entropy, or other measures of complexity in the explanation map $\Phi(\bm{x})$. The category $\Psi_{\mathrm{CMP}}$ is defined as:
$\Psi_{\mathrm{CMP}}(\Phi,f,\bm{x},C)=C\left(\Phi(\bm{x})\right),$ (35)
where $C$ represents a complexity measure, chosen based on the context of interpretability. To evaluate the complexity of explanations for RS image scene classification, we use i) Sparseness (SP)[33] and ii) Complexity (CO)[34].
@@
Sparseness (SP)[33] is a method for evaluating the sparsity of explanations and is defined as the Gini index of the explanation. It is calculated by summing the product of the ranks of the input features and their attributions and dividing by the sum of the attributions. The SP metric $\Psi_{\mathrm{SP}}$ is defined as:
@@
Complexity (CO)[34] is defined using the Shannon entropy calculation, which measures the amount of uncertainty or randomness in the attribution map. It is calculated by summing the product of the probabilities of the attributions and the logarithm of the probabilities of the attributions. The CO metric $\Psi_{\mathrm{CO}}$ is defined as:
@@
Figure 5: Explanations for a DG-ML image labeled as Forest and their corresponding SP metric scores. a) Original image; b) Pixel-wise reference map (green: Forest); c) Occlusion explanation ($\Psi_{\mathrm{SP}}=0.08$); d) LRP explanation ($\Psi_{\mathrm{SP}}=0.47$).
III D 2 Limitations
Explanation metrics in the complexity category generally assume that an explanation is simpler if it attributes relevance to fewer features for a given prediction. This assumption originates from explaining predictions for natural images, where a class is often defined by a few unique features. There, only a small fraction of the input features contribute to the final prediction. However, classification of RS images often relies more on spectral and texture features, a larger set of repetitive features is relevant, which is overlooked by complexity metrics.
@@ III D 2 Limitations
A key limitation of the Gini-based SP metric $\Psi_{\mathrm{SP}}$ is its tendency to penalize explanations with uniformly distributed attributions. In a scenario where a class spans nearly the entire image, a comprehensive explanation assigns relevance to multiple regions exhibiting discriminative spectral or textural signals. This uniform distribution is interpreted by $\Psi_{\mathrm{SP}}$ as high complexity, thereby misclassifying an accurate explanation as less concise. In contrast, an explanation for a minor class confined to a single region is favored with a low complexity score, even though such concentration does not necessarily improve interpretability. This limitation is illustrated in Fig. 5, which shows two explanations for a scene labeled as forest, a class that spans nearly the entire image. The attribution map from the Occlusion method (Fig. 5(c)) uniformly highlights all forest pixels, yielding a low SP value ($\Psi_{\mathrm{SP}}=0.08$). Although most of the input pixels are attributed as relevant, the map remains easy to interpret. By contrast, the LRP map (Fig. 5(d)) concentrates relevance on a small, localized area, producing a higher sparsity score ($\Psi_{\mathrm{SP}}=0.47$). Here the SP metric favors a compact but misleading attribution, mistaking spatial concentration for interpretive simplicity.
III E Randomization
III E 1 Background
@@
The fifth category we consider is randomization, which tests the reliability of explanations by introducing randomness into model parameters or prediction targets. These metrics evaluate the consistency of explanations under such perturbations to ensure that explanations are not artifacts of specific parameter settings or configurations. The category $\Psi_{\mathrm{RAND}}$ is defined as:
$\Psi_{\mathrm{RAND}}(\Phi,f,\bm{x},\mathcal{R},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\Phi(\bm{x},f,\mathcal{R})\right),$ (39)
where $\mathcal{R}$ represents a randomization operation (e.g., model parameter shuffling or target randomization) and SIM measures the similarity between the original explanation $\Phi(\bm{x},f)$ and the explanation generated under the randomized condition $\Phi(\bm{x},f,\mathcal{R})$. To evaluate the randomization robustness of explanation methods for RS image scene classification, we use i) Efficient Model Parameter Randomization Test (MPRT)[57], which measures explanation complexity under model parameter randomization and ii) Random Logit (RL)[36], which evaluates stability by comparing explanations for target and non-target classes using Structural Similarity Index Measure (SSIM).
@@
Efficient Model Parameter Randomization Test (MPRT)[57] measures the relative rise in the complexity of the explanation $\Phi\bigl{(}\bm{x},f_{\mathcal{R}}\bigr{)}$ from a fully randomized model $f_{\mathcal{R}}$ for image $\bm{x}$. The MPRT metric $\Psi_{MPRT}$ is defined as:
@@
Here, $\xi:\mathbb{R}^{D}\!\to\!\mathbb{R}$ denotes a complexity function. To define $\xi$, a histogram‐entropy measure grounded in Shannon entropy is employed. Specifically, the vectorized attributions $\bm{e}=\{\,a_{j,k}\}_{j,k}$ are placed into $B$ distinct bins; for the $d$-th bin, the frequency $\eta_{d}$ is computed and the corresponding normalized probability $\pi_{d}$ is
@@
The Random Logit (RL)metric [36] is defined as the SSIM over the explanation of the reference label and an explanation of non-target class $\bm{y}^{\prime}$. The RS metric $\Psi_{\mathrm{RL}}$ is defined as:
@@
where $\Phi(\bm{x},f,\hat{\bm{y}})$ is the explanation generated for the prediction $\hat{\bm{y}}$ and $\Phi\left(\bm{x},f,\bm{y}^{\prime}\right)$ is the explanation generated for a non-target class $\bm{y}^{\prime}$. Lower values indicate that the explanations are not correlated, which is desirable.
III E 2 Limitations
@@ IV A Dataset Description
BEN-v2.0 $250.249$ $19$ $2.95$ $10$ 120$\times$120 (10 m) Thematic Product
DG-ML $18.185$ $6$ $1.71$ $3$ 120$\times$120 (0.5 m) Manually Annotated
FBP-ML $109.200$ $24$ $2.61$ $4$ 256$\times$256 (4 m) Manually Annotated
BEN-v2.0 [58] is a multi-label dataset which extends the BigEarthNet archive [62] and covers ten countries in Europe. We use the Sentinel-2 imagery with 12 bands and 10 m spatial resolution. Each $1200$ m×$\times 1200$ m tile is subdivided into $120\times 120$-pixel patches. Class annotations are taken from the CORINE Land-Cover Map inventory 2018, assigning subsets of 19 LULC classes to 549 488 patches. We adopt the official train–validation–test split defined in [58].
DG-ML is a multi-label dataset, which Burgert et al. [60] derived from the original DeepGlobe Land-Cover Classification Challenge [59]. The original DeepGlobe corpus comprises 1,949 manually annotated RGB tiles of size $2448\times 2448$ pixels at 0.5 m spatial resolution, collected over Thailand, Indonesia, and India. The LULC classes are urban land, agricultural land, range land, forest, water, barren land, and unknown. Each tile was divided into a grid of $120\times 120$-pixel patches, and multi-labels were assigned from the corresponding pixel-level reference maps, while patches containing the class “unknown” were discarded. Only 20 % of single-class patches and all multi-class patches are retained, yielding 30 443 samples split into training (60%), validation (20%) and test (20%) subsets.
FBP-ML is a multi-label dataset derived from the Five Billion Pixels (FBP) collection [61], comprising 150 Gaofen-2 RGB–NIR scenes (4 m spatial resolution) that together cover roughly 50 000 km2 in China. Following the procedure used for DG-ML, each scene is partitioned into a grid of $256\times 256$-pixel patches and labeled with a 24-class scheme that distinguishes artificial, agricultural, and natural surfaces. The final dataset consists of 87 360 training samples (80 %), 10 920 validation samples (10 %), and 10 920 test samples (10 %).
IV B Evaluation Protocol
@@
The MetaQuantus framework [24] considers that the quality of an explanation cannot be objectively measured due to the lack of ground truth. It tests the explanation metrics against two failure modes: Noise Resilience (NR)of a metric, which is the resilience against minor perturbations, and Adversary Reactivity (AR), which is the reactivity of the metrics to disruptive perturbations. A perturbation is considered minor when the predicted class label does not change and disruptive when the predicted class label changes. To test against these failure modes, two consistency criteria Intra-Consistency (IAC)and Inter-Consistency (IEC)are utilized.
@@
with $0\leq\text{IEC}\leq 1$. The entries $U_{i,j}^{t}$ capture how consistently the ranking of the metrics agrees between unperturbed $\bar{\mathbf{Q}}$ and perturbed $\overline{\mathbf{Q}}^{\prime}$. For minor perturbations ($t=M$) the ordering of the $L$ methods for each sample must remain identical, whereas for disruptive perturbations ($t=D$) the unperturbed score must always ranked higher than its perturbed counterpart.
@@
with the all-ones vector representing an ideal explanation metric that achieves a value of 1 on all criteria. A metric with an ${MC}$ score close to 1 indicates strong performance across all criteria tested with higher values reflecting greater overall robustness and reliability. The MC score is averaged over two types of perturbation: model perturbation and input perturbation. In model perturbation, random noise is added to the parameters, whereas in input perturbation, it is added to the input data.
IV C Experimental Setup
@@
We evaluated the quality of the explanation using the considered metrics and evaluated their reliability with MetaQuantus [24] on 512 randomly drawn samples, a sample size demonstrated to be sufficient in previous work [24]. The hyperparameter were set as in the original publication ($k=5$, three iterations). Gaussian noise was chosen for the input perturbation to preserve physical consistency, in line with the recommendations in [63]. The choice of model depth does not substantially affect the outcome of the evaluation, as the performance of MetaQuantus has been shown to be largely invariant with respect to model architecture and depth [24]. For model training, we used the AdamW optimizer, initialized at a learning rate of $10^{-2}$ and a weight decay of $0.9$. All experiments were performed on Nvidia H100 80GB GPUs.
@@
We limit our analysis to five explanation methods, as prior work [24] indicates a minimal influence of the number of methods on the result scores. Additionally, we used a Random baseline explanation method that samples values randomly from a uniform distribution $U(0,1)$. For a multi-label case with a label set $\{1,…,K\}$, we apply the same method separately to each label $c\in\{1,…,K\}$ to obtain a feature attribution map for each class $c$. Furthermore, as common in the literature, we summed the feature attributions over the channels.
@@
The hyperparameters for the explanation methods were configured as follows. For DeepLIFT, the perturbation baseline value was set to 00. For GradCAM, the last convolutional layer was analyzed and the explanations were upscaled by bilinear interpolation. The LIME method employed a Euclidean exponential kernel (kernel width 500) as a similarity measure, used linear regression as an interpretable model, and generated 15 superpixels using SLIC. For LRP, following [45], the first third of the network was initialized with the LRP-$\gamma$ rule, the second third with the LRP-$\epsilon$ rule and the final third with the LRP-00 rule. Lastly, the Occlusion method employed a sliding window of size $(50,50)$ with strides of $(10,10)$ across the channels and the perturbation baseline was set to 00. The quality of the explanations was assessed using 10 explanation metrics implemented via the Quantus framework [25] with default hyperparameters. Quantus is recognized as the most suitable framework for evaluating explanations [64]. The explanation metrics were computed over 1024 randomly drawn test samples from each dataset. The explanation methods are implemented using the Captum framework [65] with modifications to Captum, Quantus, and MetaQuantus to support multi-label image scene classification.

### v3

@@ III.C.1 Background
If all top-$K$ pixels are within the reference mask $S_{gt}$, then $\Psi_{\mathrm{TKI}}=1$, therefore, a higher score is desired.
@@
The RRA metric [31] measures the concentration of high-intensity relevance values within the pixel-level reference map. Let $|S_{\text{tot}}|$ be the size of the reference mask. Let us assume that $K$ pixels with the highest relevance values are selected from $\Phi(\bm{x})$. We then count how many of these selected pixels lie within the reference pixels and divide this number by $|S_{\text{tot}}|$:
$\Psi_{\mathrm{RRA}}=\frac{|P_{\text{top }K}\cap S_{\text{tot}}|}{|S_{\text{tot}}|}.$ (24)
A higher $\Psi_{\mathrm{RRA}}$ indicates stronger localization, as more highly relevant pixels fall within the reference region.
III.C.2 Limitations
In RS scene classification, label dependencies arise when the presence of one label influences the likelihood of another, e.g., ships typically co-occur with sea. Localization metrics do not account for such dependencies. Their reliability also depends on accurate pixel-level reference maps, which are costly to obtain and prone to errors. Outdated or inaccurate annotations compromise the validity of these metrics.
@@ III.C.2 Limitations
Additionally, most localization metrics become unreliable when the target class spans a large area of the input image. Let us consider the TKI metric in the scenario where we have a single label for the input image. Here, the binary reference mask $\mathbf{s}_{gt}$ for the target label becomes a matrix of ones, encompassing all pixels. Thus, $\Psi_{\mathrm{TK}}$ inherently achieves its maximum value of 1, regardless of the explanation $\bm{e}$. Since $\mathbf{s}_{gt}$ includes every pixel, the intersection $\mathbf{s}_{1-K}=\mathbf{r}_{1-K}\cap\mathbf{s}_{gt}$ reduces to $\mathbf{r}_{1-K}$ itself, as all top-ranked indices in $\mathbf{r}_{1-K}$ necessarily lie within $\mathbf{s}_{gt}$. Consequently, the numerator and denominator of $\Psi_{\mathrm{TK}}$ both equal $K$, yielding $\Psi_{\mathrm{TK}}=1$. In this case, the metric cannot differentiate meaningful explanations from arbitrary ones. This phenomenon is well-documented for simpler localization metrics [57], such as the Pointing Game [58]. However, in RS scene classification, this also creates challenges for more sophisticated metrics, as in most RS scene classification tasks, as LULC classes tend to span larger areas.
III.D Complexity
III.D.1 Background
@@
The fourth category we consider is complexity, which quantifies the simplicity and interpretability of explanations, ensuring that they are useful for human understanding. These metrics evaluate properties such as sparsity, entropy, or other measures of complexity in the explanation map $\Phi(\bm{x})$. To evaluate the complexity of explanations for RS image scene classification, we use i) Sparseness (SP) [33]; and ii) Complexity (CO) [34].
@@
SP [33] is a method for evaluating the sparsity of explanations and is defined as the Gini index of the explanation. It is calculated by summing the product of the ranks of the input features and their attributions and dividing by the sum of the attributions. The SP metric $\Psi_{\mathrm{SP}}$ is defined as:
@@
CO [34] is defined using the Shannon entropy calculation, which measures the amount of uncertainty or randomness in the attribution map. It is calculated by summing the product of the probabilities of the attributions and the logarithm of the probabilities of the attributions. The CO metric $\Psi_{\mathrm{CO}}$ is defined as:
@@
Figure 5: Explanations for an image from the DG-ML dataset [46] labeled as Forest and their corresponding SP metric scores. a) Original image; b) Pixel-wise reference map (green: Forest); c) Occlusion explanation ($\Psi_{\mathrm{SP}}=0.08$); d) LRP explanation ($\Psi_{\mathrm{SP}}=0.47$).
III.D.2 Limitations
@@
Explanation metrics in the complexity category generally assume that an explanation is simpler if it attributes relevance to fewer features for a given prediction. This assumption originates from natural images, where classes are defined by a few distinctive features, and only a small fraction of pixels contribute to the prediction. However, as discussed before, in RS scenes LULC classes often cover large areas with repeated class-relevant features, so relevance is distributed across many pixels. Complexity metrics can therefore underestimate explanation quality. A key limitation of the Gini-based SP metric $\Psi_{\mathrm{SP}}$ is its tendency to penalize explanations with uniformly distributed attributions. In a scenario where a class spans nearly the entire image, a comprehensive explanation assigns relevance to multiple regions exhibiting discriminative spectral or textural signals. This uniform distribution is interpreted by $\Psi_{\mathrm{SP}}$ as high complexity, thereby misclassifying an accurate explanation as less concise. In contrast, an explanation for a minor class confined to a single region is favored with a low complexity score, even though such concentration does not necessarily improve interpretability. This limitation is illustrated in Fig. 5, which shows two explanations for an image labeled as forest, a class that spans nearly the entire image. The attribution map from the Occlusion method (Fig. 5(c)) uniformly highlights all forest pixels, yielding a low SP value ($\Psi_{\mathrm{SP}}=0.08$). Although most of the input pixels are attributed as relevant, the map remains easy to interpret. On the contrary, the LRP map (Fig. 5(d)) concentrates relevance on a small, localized area, producing a higher sparsity score ($\Psi_{\mathrm{SP}}=0.47$). Here, the SP metric favors a compact but misleading attribution, mistaking spatial concentration for interpretive simplicity.
III.E Randomization
III.E.1 Background
@@
The fifth category we consider is randomization, which tests the reliability of explanations by introducing randomness into model parameters or prediction targets. These metrics evaluate the consistency of explanations under such perturbations to ensure that explanations are not artifacts of specific parameter settings or configurations. Here, we use i) Efficient Model Parameter Randomization Test (MPRT) [59], which measures explanation complexity under model parameter randomization, and ii) Random Logit (RL) [36], which evaluates stability by comparing explanations for target and non-target classes using Structural Similarity Index Measure (SSIM).
@@
MPRT [59] measures the relative increase in the complexity of the explanation $\Phi\bigl(\bm{x},f_{\mathcal{R}}\bigr)$ from a fully randomized model $f_{\mathcal{R}}$ for image $\bm{x}$. The MPRT metric $\Psi_{MPRT}$ is defined as:
@@
Here, $\xi:\mathbb{R}^{D}\!\to\!\mathbb{R}$ denotes a complexity function. To define $\xi$, a histogram‐entropy measure grounded in Shannon entropy is employed. Specifically, the vectorized attributions $\bm{e}=\{\,a_{j,k}\}_{j,k}$ are placed into $B$ distinct bins. For the $d$-th bin, the frequency $\eta_{d}$ is computed and the corresponding normalized probability $\pi_{d}$ is
@@
The RL metric [36] is defined as the SSIM over the explanation of the reference label and an explanation of the non-target class $\bm{y}^{\prime}$. The RS metric $\Psi_{\mathrm{RL}}$ is defined as:
@@
[…]
III.E.2 Limitations
@@ IV.A Dataset Description
BEN-v2.0 [60] $250.249$ 1919 $2.95$ $10$ $\times120$ (10 m) Thematic Product
DG-ML [46] $18.185$ 66 $1.71$ 33 $\times120$ (0.5 m) Manually Annotated
FBP-ML [62] $109.200$ $24$ $2.61$ 44 $\times256$ (4 m) Manually Annotated
BigEarthNet-S2 v2.0 (denoted as BEN-v2.0 hereafter) [60] is a multi-label dataset, including 549 488 Sentinel-2 images acquired over ten countries in Europe. We used all the bands of the Sentinel-2 images, and thus applied cubic interpolation to 20-m and 60-m bands. Each image is associated with one or more class labels (i.e., multi-labels) based on the 19 class nomenclature [63]. In the experiments, we use the official train-validation-test split defined in [60].
DG-ML is a multi-label dataset, which Burgert et al. [61] derived from the original DeepGlobe Land-Cover Classification Challenge dataset [46]. The original DeepGlobe corpus comprises 1,949 manually annotated RGB tiles of size $2448\times 2448$ pixels at 0.5 m spatial resolution, collected over Thailand, Indonesia, and India. Each tile is divided into 400 images, each of which is a section of $120\times 120$ pixel images. Multi-labels were assigned from the corresponding pixel-level reference maps, while images containing the class “unknown” were discarded. Only 20% of single-class images and all multi-class images are retained, yielding 30 443 samples split into training (60%), validation (20%), and test (20%) subsets.
FBP-ML is a multi-label dataset derived from the Five Billion Pixels (FBP) collection [62], comprising 150 Gaofen-2 RGB–NIR scenes (4 m spatial resolution) that together cover roughly 50 000 km2 in China. Following the procedure used for DG-ML, each tile is partitioned into an area of $256\times 256$-pixel images and labeled with a 24-class scheme that distinguishes artificial, agricultural, and natural surfaces. The final dataset consists of a training set with 87 360 samples (80%), a validation set with 10 920 samples (10%), and a test set with 10 920 samples (10%).
IV.B Evaluation Protocol
@@
As the quality of an explanation cannot be objectively measured due to the lack of ground truth [24], the explanation metrics are tested against two failure modes: Noise Resilience (NR) of a metric, which is the resilience against minor perturbations, and Adversary Reactivity (AR), which is the reactivity of the metrics to disruptive perturbations. A perturbation is considered minor when the predicted class label does not change, and disruptive when the predicted class label changes. To test against these failure modes, two consistency criteria Intra-Consistency (IAC) and Inter-Consistency (IEC) are utilized.
@@
with $0\leq\text{IEC}\leq 1$. The entries $U_{i,j}^{t}$ capture how consistently the ranking of the metrics agrees between unperturbed $\bar{\mathbf{Q}}$ and perturbed $\overline{\mathbf{Q}}^{\prime}$. For minor perturbations ($t=M$) the ordering of the $L$ methods for each sample must remain identical, whereas for disruptive perturbations ($t=D$), the unperturbed score must always be ranked higher than its perturbed counterpart.
@@
with the all-ones vector representing an ideal explanation metric that achieves a value of 1 on all criteria. A metric with an $\mathbf{MC}$ score close to 1 indicates strong performance across all criteria tested, with higher values reflecting greater overall robustness and reliability. The $\mathbf{MC}$ score is averaged over two types of perturbation: model perturbation and input perturbation. In model perturbation, random noise is added to the parameters, whereas in input perturbation, it is added to the input data.
While the protocol provides a systematic assessment of explanation metrics, it is computationally demanding as its runtime depends on the efficiency of the employed explanation methods and metrics. In particular, perturbation-based methods and metrics require multiple forward passes and repeated perturbations, which significantly increases computation time compared to closed-form or gradient-based approaches. This limitation is not inherent to the evaluation protocol but reflects the methodological design of certain metrics. A detailed comparison of computational efficiency across methods and metrics is provided in subsection V.C.
IV.C Experimental Setup
@@
We evaluated the quality of the explanation using the considered metrics and evaluated their reliability with MetaQuantus [24] on 512 randomly drawn samples, a sample size demonstrated to be sufficient in previous work [24]. The hyperparameters were set as in the original publication ($k=5$, three iterations, which were shown in [24] to be effective across different datasets. Gaussian noise was chosen for the input perturbation to preserve physical consistency, in line with the recommendations in [64]. The choice of model depth does not substantially affect the outcome of the evaluation, as the performance of MetaQuantus has been shown to be largely invariant with respect to model architecture and depth [24]. For model training, we used the AdamW optimizer, initialized at a learning rate of $10^{-2}$ and a weight decay of $0.9$. All experiments were performed on Nvidia H100 80GB GPUs.
@@
We limit our analysis to five explanation methods, as prior work [24] indicates a minimal influence of the number of methods on the result scores. Additionally, we used a Random baseline explanation method that samples values randomly from a uniform distribution $U(0,1)$. For a multi-label case with a label set $\{1,…,K\}$, we apply the same method separately to each label $c\in\{1,…,K\}$ to obtain a feature attribution map for each class $c$. Furthermore, as is common in the literature, we summed the feature attributions over the channels.
@@
The hyperparameters for the explanation methods were configured as follows. For DeepLIFT, the perturbation baseline value was set to 0. For GradCAM, the last convolutional layer was analyzed, and the explanations were upscaled by bilinear interpolation. The LIME method employed a Euclidean exponential kernel (kernel width 500) as a similarity measure, used linear regression as an interpretable model, and generated 15 superpixels using SLIC. For LRP, following [50], the first third of the network was initialized with the LRP-$\gamma$ rule, the second third with the LRP-$\epsilon$ rule and the final third with the LRP-0 rule. Lastly, the Occlusion method employed a sliding window of size $(50,50)$ with strides of $(10,10)$ across the channels, and the perturbation baseline was set to 0. The quality of the explanations was assessed using 10 explanation metrics implemented via the Quantus framework [25] with default hyperparameters. Quantus is recognized as the most suitable framework for evaluating explanations [65]. The explanation metrics were computed over 1024 randomly drawn test samples from each dataset. Both explanation methods and metrics were used with their standard parameters to evaluate the configurations that are most commonly used in RS. The explanation methods are implemented using the Captum framework [66] with modifications to Captum, Quantus, and MetaQuantus to support multi-label image scene classification.
Figure 6: Mean $\mathbf{MC}$ scores for the considered explanation metrics: a) scores associated to all considered RS datasets (DG-ML [46], FBP-ML [62] and BEN-v2.0 [60]); b) scores associated to each individual dataset (DG-ML [46], FBP-ML [62] and BEN-v2.0 [60]). Metrics are categorized into five color-coded groups: i) Complexity (brown), ii) Faithfulness (blue), iii) Localization (orange), iv) Randomization (green), and v) Robustness (red).